OpenStudy (ikram002p):

prove or disprove if \(\lim (g(x)) =\infty\) and \(\lim(f(x)) =\infty\) then \( \lim (g(x).f(x)) =\infty \) as x goes to some value.

OpenStudy (pawanyadav):

lim(g(x).f(x))=lim(g(x)).lim(f(x))=~×~=~ Am representing infinity by ~ Hence proved.

OpenStudy (ikram002p):

so ur saying infinite*infinite=infinite which is away impossible mathematically. you can't just touch infinity lol

OpenStudy (pawanyadav):

Ha ha ....

OpenStudy (ikram002p):

unless u have a way to convince that was true.

OpenStudy (pawanyadav):

It is. Correct.

OpenStudy (pawanyadav):

You can check...

OpenStudy (ikram002p):

although i'm going now. i shall see this next morning, gn=)

OpenStudy (pawanyadav):

Wait....

OpenStudy (pawanyadav):

~×~=~ This is correct.

OpenStudy (ikram002p):

well @wolfram is not convincing enough if u think about that.

zepdrix (zepdrix):

I can't quite remember how to prove these limit rules :) But my book has a good explanation.. trying to find it >.< grr

OpenStudy (pawanyadav):

In limits......it is correct ,it don't come under 7 indeterminate forms.

OpenStudy (ikram002p):

can't wait to see @zepdrix =D

OpenStudy (ikram002p):

well why it's correct is part of my question @Pawanyadav :)

OpenStudy (pawanyadav):

I'm using that, as a result ,I can't prove that....

OpenStudy (pawanyadav):

I prove the question ,that's final....

zepdrix (zepdrix):

Grr I dunno ;c maybe I was mistaken..

OpenStudy (pawanyadav):

Is anyone saying that ~×~=~ in limits is false.

OpenStudy (irishboy123):

in limits?!?!

zepdrix (zepdrix):

No, that sounds right :) I just can't figure out how to justify it hehe

OpenStudy (irishboy123):

so what is ~ + ~ using the new nomenclature?

OpenStudy (pawanyadav):

That's what am saying, it is correct.. But @ikram002p is not ready to accept it

zepdrix (zepdrix):

He was representing \(\rm \infty\cdot\infty\) by typing ~x~

OpenStudy (pawanyadav):

@Irishboy123 ,it is already ~ About which new nomenclature are you talking

OpenStudy (pawanyadav):

~+~=~ ~×~=~ But. ~÷~ and ~-~ are indeterminate forms. ..do anyone have any problem with this.

OpenStudy (pawanyadav):

@ikram002p as it is proved ,don't forget the bet now

OpenStudy (thomas5267):

Both f(x) and g(x) blow up to infinity with finite but possibly different x right? If both f(x) and g(x) only blow up at x to infinity, then take f(x)=g(x)=x. f(x).g(x)=x^2 and only blows up at infinity, which seems to contradict your requirement that f(x).g(x) blows up at some, presumably finite, value of x.

OpenStudy (thomas5267):

Do tan(x) and cot(x) satisfy your reqirement? Both blow up at different but finite x. The product of two is 1 and does not blow up.

ganeshie8 (ganeshie8):

Both the limits are infinity. So, given any large positive number \(\epsilon\), we can find \(L\) and \(M\) such that \(f(x)\gt \epsilon\) and \(g(x)\gt \epsilon\) for all \(x\gt \max\{L,M\}\). It follows \(f(x)*g(x)\gt \epsilon\) whenever \(x\gt \max\{L,M\}\). That is \(f(x)*g(x)\) arbitrarily increases for large values of \(x\). Same as saying \(\lim f(x)*g(x) = \infty \).

OpenStudy (pawanyadav):

@thomas5267. here is only 0ne x not two different x both f(x) and g(x) tends to infinity at some but same value of x.

OpenStudy (ikram002p):

=)